If Marcelle deposits $250,000 into a mutual fund which pay an annual interest rate of 5% compounded continuously. Assuming that there are no further deposits and there were no withdrawals, how long must she wait before her investment reaches $1 million?
Answer by stanbon(75887)

If Marcelle deposits $250,000 into a mutual fund which pay an annual interest rate of 5% compounded continuously. Assuming that there are no further deposits and there were no withdrawals, how long must she wait before her investment reaches $1 million?
----------------------------
A(t) = Ao*e^(rt)
1,000,000 = 250,000*e^(0.05t)
e^(0.05t) = 4
----
Take the natural log of both sides:
0.05t = 4
---
t = 80 years
=====================
Cheers,
Stan H.
===============